• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name fansler

    Log entry time 13:52:37 on July 20, 2010

    Entry number 331635

    keyword=Left arm magnet temps

    The left arm Q3 temps were not updating. B Carlton suggested we reseat the cards. Slot 17 and 18 appeared to be locked up, we reseated the cards and they are updating again. We checked the right and found slot 18 not functioning. Reseating the card did not bring it back online, so we will have to replace it.